OpenStudy (anonymous):

Find \((-1+\sqrt{2} i)^3\) using DeMoivre's Theorem.

OpenStudy (anonymous):

first job is to write this in polar (trig) form do you know how to do that?

OpenStudy (anonymous):

yea i got\[\sqrt{3}^3(cos(-0.955)+i~sin(-0.955))\]butI dont know if im supposed to add \(\pi\) or \(2\pi\)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

ok lets back up a second

OpenStudy (anonymous):

\(r\left(\cos(\theta)+i\sin(\theta)\right)\) is the form you want, and so you need \(r\) and \(\theta\)

OpenStudy (anonymous):

\(r\) is easiest. if you have \(a+bi\) then \(r=\sqrt{a^2+b^2}\)

OpenStudy (anonymous):

\[r=\sqrt{(-1)^2+(\sqrt{2})^2}\]\[r=\sqrt{3}\]\[\theta=tan^{-1}\left({\sqrt{2}\over-1}\right)\]\[\theta\approx-0.955\]

OpenStudy (anonymous):

in your case you have \(-1+\sqrt{2}i\) and so \(r=\sqrt{(-1)^2+\sqrt{2}^2}=\sqrt{1+2}=\sqrt{3}\)

OpenStudy (anonymous):

ok good for \(r\) lose the calculator for \(\theta\) it is not helping you

OpenStudy (anonymous):

if you were in quadrant 1, then you can use \(\theta=\tan^{-1}(\frac{b}{a})\) but you are in quadrant 3

OpenStudy (anonymous):

how do i know that im in quadrant 3?

OpenStudy (anonymous):



OpenStudy (anonymous):

that's quadrant 2..

OpenStudy (anonymous):

i lied, you are in quadrant 2

OpenStudy (anonymous):

but in any case \(\tan^{-1}(x)\) only gives answers in quad 1 or 4

OpenStudy (anonymous):

right so i would add \(\pi\) to to \(-0.955\) and get \(2.186\) then

OpenStudy (anonymous):

which is in quadrant 2

OpenStudy (anonymous):

yeah i guess that is what you need, right

OpenStudy (anonymous):

now you can multiply that angle by 3, and cube \(\sqrt{3}\) to get your answer

OpenStudy (anonymous):

just for my own information, you are sure this is not \(-1+\sqrt{3}i\) right? that would be much easier

OpenStudy (anonymous):

when you multiply that by 3 .. you get 6.559 which means i would subtract \(2\pi\) to get the angle within the range \(0\le\theta\le2\pi\) and that would be \(0.276\)

OpenStudy (anonymous):

no it is what i posted, though it would be easier if it was so is that right?

OpenStudy (anonymous):

@satellite73